# Text-to-Image Generation Via Energy-Based CLIP

### v1

@@ Appendix C Training protocol
In algorithm 1, we detail to overall training procedure of CLIP-JEM. First, we draw image-text batches from dataset $\mathcal{D}$ and apply $T_{\text{adv}}$ iterations to obtain the adversarial visual inpputs $\mathbf{I}_{\text{adv}}$ which we use to calculate the adversarial loss $\mathcal{L}_{\text{adv}}$. Next, we perform an iterative pixel-space optimization to obtain “negative” samples $\tilde{\mathbf{I}}$. Specifically, this is done using $T_{\text{JEM}}$ iterations using AdamW optimizer. We form our joint energy based loss using the “positive” and “negative” samples. Lastly, we combine these two terms to formulate our overall objective, which we use to update the vision encoder weights. We repeat this process until convergence.
@@
       Sample $(\mathbf{I},\mathbf{T})$ from dataset $\mathcal{D}$ Contrastive adversarial loss $\mathbf{\delta}_{0}\leftarrow\mathbf{0}$ for t from 0 to $T_{\text{adv}}$ do
@@
      $\mathbf{I}_{\text{adv}}=\mathbf{I}+\mathbf{\delta}_{T_{\text{adv}}}$ $\mathcal{L}_{\text{adv}}=\operatorname{ClipLoss}(f_{\theta}^{I}(\mathbf{I}_{adv}),f_{\theta}^{T}(\mathbf{T}))$ Contrastive energy loss Sample initial “negative” sample $\tilde{\mathbf{I}}$. $\text{Optimizer}\leftarrow\text{AdamW}(\text{params}=\tilde{\mathbf{I}},\text{lr}=\alpha_{2})$ for t from 0 to $T_{\text{JEM}}$ do
             $\mathcal{L}_{\text{JEM}}=\operatorname{ClipLoss}(f_{\theta}^{I}(\tilde{\mathbf{I}}+\beta\mathbf{n}),f_{\theta}^{T}(\mathbf{T}))$ $\mathbf{n}\sim\mathcal{N}(0,\mathbf{I}),$$\beta$ is a small scalar Calculate $\nicefrac{{\partial{\mathcal{L}_{\text{JEM}}}}}{{\partial{\tilde{\mathbf{I}}}}}$ and perform an optimizer step
@@
      $\mathcal{L}_{\text{JEM}}=\operatorname{ClipLoss}(f_{\theta}^{I}(\operatorname{Concat}(\mathbf{I},\tilde{\mathbf{I}})),f_{\theta}^{T}(\mathbf{T}))$ Update the vision encoder $\mathcal{L}=\mathcal{L}_{\text{adv}}+\gamma*\mathcal{L}_{\text{JEM}}$ Calculate $\nicefrac{{\partial{\mathcal{L}}}}{{\partial{\theta}}}$ and update CLIP image encoder $f_{\theta}^{I}(\cdot)$
@@ Appendix D Ablation study
In table 1, we report the results of both ViT-B-32 and ConvNext base using CLIP-JEM. Notably, despite the two architectures share a similar capacity, the ConvNext FID score is significantly better (by $\mathbf{33.5}$ points). We hypothesize that this stems from the improved prior that CNN-based architectures serves. Additionally, generated images from ViT contain grid artifact from the patch processing mechanism. Thus, we mainly focus on ConvNext-based models.
@@
To highlight the importance of combining the adversarial and energy contrastive losses, we train the same model for $1,000$ iterations using (i) contrastive energy loss; (ii) contrastive adversarial loss; and (iii) contrastive energy loss + contrastive adversarial loss. We plot the results of $16$ text prompts from MS-COCO for the resulting models in fig. 8. The results indicate that using only the contrastive energy loss does not produce meaningful outputs. In contrast, employing solely the contrastive adversarial loss results in meaningful but unrealistic content. Remarkably, combining the two objectives (CLIP-JEM’s approach) leverages the strengths of both CLIPAG and JEMs, yielding superior outcomes.
Figure 8: Objectives ablation study. CLIP-JEM employs a two objectives: Contrastive Adversarial Loss and a Contrastive Energy Loss. In this figure we ablate the effect of each objective, highlighting the importance of each objective. We plot $16$ images based on different textual prompts from MS-COCO after $500$ training steps.
@@ Appendix E Connection to Stochastic Gradient Langevin Dynamics
where $\alpha$ is the step size and $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\alpha\bm{I})$ is a Gaussian noise. From eq. 6, the above equation can be also written in terms of energy,
@@
Note that in our algorithm we add a small Gaussian perturbation to the input, prior to the calculation of the gradient (algorithm 1). This introduces a similar, yet different, stochasticity in our approach, compared with SGLD. To better understand this, we expand the following Taylor series:
@@
We demonstrate the randomness introduced by this mechanism using our sampling process in fig. 9. Unlike to SGLD, we utilize an AdamW optimizer for the update step for the generated sample. Thus, we do not use the raw gradient as in SGLD since our optimizer has an adaptive learning rate mechanism. This results in a different effective step size per iteration ($\alpha$ in SGLD). AdamW also enables momentum term, which introduces bias to the gradients. During CLIP-JEM training, we utilize the momentum as it significantly stabilizes the training. However, at inference time, we do not employ momentum, resulting in an unbiased version of the gradients.
@@ Appendix F Additional results
We provide additional qualitative results for both improving CLIP-based generative frameworks and text-to-image generation. For the former, we provide the results using the same prompts with augmentations (figs. 11 and 12). As can be seen, seamlessly replacing CLIP with CLIP-JEM leads to improved results with and without augmentations. Notably, CLIP-JEM results without augmentations are comparable to the ones of CLIP with augmentation, offering a substantial reduce of computational overhead. For the latter, we provide more generated image (fig. 10) and demonstrate the stochasticity of CLIP-JEM(fig. 9).

### v2

@@ Appendix C Training protocol
In algorithm˜1, we detail to overall training procedure of CLIP-JEM. First, we draw image-text batches from dataset $\mathcal{D}$ and apply $T_{\text{adv}}$ iterations to obtain the adversarial visual inpputs $\mathbf{I}_{\text{adv}}$ which we use to calculate the adversarial loss $\mathcal{L}_{\text{adv}}$. Next, we perform an iterative pixel-space optimization to obtain “negative” samples $\tilde{\mathbf{I}}$. Specifically, this is done using $T_{\text{JEM}}$ iterations using AdamW optimizer. We form our joint energy based loss using the “positive” and “negative” samples. Lastly, we combine these two terms to formulate our overall objective, which we use to update the vision encoder weights. We repeat this process until convergence.
@@
   Sample $(\mathbf{I},\mathbf{T})$ from dataset $\mathcal{D}$ /* Contrastive adversarial loss */ $\mathbf{\delta}_{0}\leftarrow\mathbf{0}$ for t from 0 to $T_{\text{adv}}$ do
@@
 $\mathbf{I}_{\text{adv}}=\mathbf{I}+\mathbf{\delta}_{T_{\text{adv}}}$ $\mathcal{L}_{\text{adv}}=\operatorname{ClipLoss}(f_{\theta}^{I}(\mathbf{I}_{adv}),f_{\theta}^{T}(\mathbf{T}))$ /* Contrastive energy loss */ Sample initial “negative” sample $\tilde{\mathbf{I}}$. $\text{Optimizer}\leftarrow\text{AdamW}(\text{params}=\tilde{\mathbf{I}},\text{lr}=\alpha_{2})$ for t from 0 to $T_{\text{JEM}}$ do
    $\mathcal{L}_{\text{JEM}}=\operatorname{ClipLoss}(f_{\theta}^{I}(\tilde{\mathbf{I}}+\beta\mathbf{n}),f_{\theta}^{T}(\mathbf{T}))$ /* $\mathbf{n}\sim\mathcal{N}(0,\mathbf{I}),$$\beta$ is a small scalar */ Calculate $\nicefrac{{\partial{\mathcal{L}_{\text{JEM}}}}}{{\partial{\tilde{\mathbf{I}}}}}$ and perform an optimizer step
@@
 $\mathcal{L}_{\text{JEM}}=\operatorname{ClipLoss}(f_{\theta}^{I}(\operatorname{Concat}(\mathbf{I},\tilde{\mathbf{I}})),f_{\theta}^{T}(\mathbf{T}))$ /* Update the vision encoder */ $\mathcal{L}=\mathcal{L}_{\text{adv}}+\gamma*\mathcal{L}_{\text{JEM}}$ Calculate $\nicefrac{{\partial{\mathcal{L}}}}{{\partial{\theta}}}$ and update CLIP image encoder $f_{\theta}^{I}(\cdot)$
@@ Appendix D Ablation study
In table˜1, we report the results of both ViT-B-32 and ConvNext base using CLIP-JEM. Notably, despite the two architectures share a similar capacity, the ConvNext FID score is significantly better (by $\mathbf{33.5}$ points). We hypothesize that this stems from the improved prior that CNN-based architectures serves. Additionally, generated images from ViT contain grid artifact from the patch processing mechanism. Thus, we mainly focus on ConvNext-based models.
@@
To highlight the importance of combining the adversarial and energy contrastive losses, we train the same model for 1,0001,0001 , 000 iterations using (i) contrastive energy loss; (ii) contrastive adversarial loss; and (iii) contrastive energy loss + contrastive adversarial loss. We plot the results of 161616 text prompts from MS-COCO for the resulting models in fig.˜7. The results indicate that using only the contrastive energy loss does not produce meaningful outputs. In contrast, employing solely the contrastive adversarial loss results in meaningful but unrealistic content. Remarkably, combining the two objectives (CLIP-JEM’s approach) leverages the strengths of both CLIPAG and JEMs, yielding superior outcomes.
Figure 7: Objectives ablation study. CLIP-JEM employs a two objectives: Contrastive Adversarial Loss and a Contrastive Energy Loss. In this figure we ablate the effect of each objective, highlighting the importance of each objective. We plot 161616 images based on different textual prompts from MS-COCO after 500500500 training steps.
@@ Appendix E Connection to Stochastic Gradient Langevin Dynamics
where $\alpha$ is the step size and $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\alpha\bm{I})$ is a Gaussian noise. From eq.˜7, the above equation can be also written in terms of energy,
@@
Note that in our algorithm we add a small Gaussian perturbation to the input, prior to the calculation of the gradient (algorithm˜1). This introduces a similar, yet different, stochasticity in our approach, compared with SGLD. To better understand this, we expand the following Taylor series:
@@
We demonstrate the randomness introduced by this mechanism using our sampling process in fig.˜8. Unlike to SGLD, we utilize an AdamW optimizer for the update step for the generated sample. Thus, we do not use the raw gradient as in SGLD since our optimizer has an adaptive learning rate mechanism. This results in a different effective step size per iteration ($\alpha$ in SGLD). AdamW also enables momentum term, which introduces bias to the gradients. During CLIP-JEM training, we utilize the momentum as it significantly stabilizes the training. However, at inference time, we do not employ momentum, resulting in an unbiased version of the gradients.
@@ Appendix F Additional results
We provide additional qualitative results for both improving CLIP-based generative frameworks and text-to-image generation. For the former, we provide the results using the same prompts with augmentations (figs.˜10 and 11). As can be seen, seamlessly replacing CLIP with CLIP-JEM leads to improved results with and without augmentations. Notably, CLIP-JEM results without augmentations are comparable to the ones of CLIP with augmentation, offering a substantial reduce of computational overhead. For the latter, we provide more generated image (fig.˜9) and demonstrate the stochasticity of CLIP-JEM(fig.˜8).
Appendix G Positioning CLIP-JEM within the Landscape of Generative Models
Comparison with Diffusion Models
While CLIP-JEM does not achieve state-of-the-art FID scores in text-to-image generation (see table˜1), it offers several notable advantages. It exhibits superior compositional generalization, outperforming Stable Diffusion v1.4 and v2 on the T2I-CompBench benchmark (Gadre et al., 2024), as well as surpassing compositionality-oriented diffusion-based methods. This performance stems from its joint energy–contrastive adversarial training, which enhances the alignment between textual prompts and generated images. Unlike diffusion pipelines that rely on multi-step denoising in pixel or latent spaces - necessitating a learned decoder and intricate noise scheduling - CLIP-JEM samples directly in pixel space through a straightforward iterative gradient-following process. This approach yields energy values proportional to sample likelihoods, facilitating explicit probabilistic ranking and selection of outputs. We leverage this property to develop a robust text-to-image alignment metric (see section˜4.3) and to convert unconditional diffusion models into text-conditioned generators. Furthermore, CLIP-JEM ’s energy-based formulation allows for seamless "plug-and-play" integration into CLIP-based synthesis workflows; for example, replacing the CLIP encoder in CLIPDraw with CLIP-JEM results in more semantically faithful renderings under the same optimization loop (Frans et al., 2022).
Comparison with Architectures Employing CLIP Image Encoders
Leading text-to-image systems incorporate CLIP’s image encoder alongside distinct generative backbones rather than utilizing it to generate pixels directly. DALL·E 2 (Ramesh et al., 2022) employs a diffusion prior to map CLIP text embeddings to CLIP image embeddings, followed by a modified GLIDE decoder for pixel synthesis. DiffusionCLIP (Kim et al., 2022) integrates a CLIP-based contrastive loss into the reverse diffusion trajectory of pretrained diffusion models to enable zero-shot, text-guided image manipulation. VQGAN-CLIP (Crowson et al., 2022) pairs a pretrained VQGAN generator with CLIP’s image encoder to score and steer pixel-space updates via gradient descent. Recently, BLIP3-o (Chen et al., 2025) utilizes the CLIP image encoder in conjunction with a large language model and a diffusion transformer to generate images. In contrast, CLIP-JEM constructs a generative model directly from CLIP itself, eliminating the need for an additional generative model.
Appendix H Limitation and Futrue Work
Our current work exhibits two main limitations. First, CLIP-JEM shows limited diversity in its generated samples. The primary source of stochasticity comes from the random initialization process, and while we experimented with noise injection during optimization steps, we maintained a small noise coefficient for stability. This resulted in generated images sharing similar visual characteristics for the same prompt, as evident in fig.˜8. Future work could explore the enhancement of sample diversity through increased noise levels during optimization or alternative initialization strategies. Second, CLIP-JEM underperforms on tasks involving spatial relationships, as demonstrated in table˜2. This limitation stems from CLIP’s inherent constraints in spatial compositionality, as our model initializes from CLIP weights. This aligns with the well-documented limitations of CLIP in handling spatial relationships. Future research could focus on addressing these spatial understanding capabilities, potentially through architectural modifications or specialized training objectives that better capture spatial information.